OpenStudy (anonymous):

To determine the distance between two points A and B, a surveyor chooses a point C that is 295 yards from A and 225 yards from B. If ∠BAC has a measure of 49.5°, approximate the distance between A and B. (Round your answer to the nearest whole number.)

OpenStudy (imstuck):

You need the Law of Sines here. Do you know it?

OpenStudy (anonymous):

\[\frac{ a }{\sin A }+\frac{ b }{ \sin B }= \frac{ c }{ \sin C}\]

OpenStudy (imstuck):

Well, it does not have a + sign in there. Only = signs between all of them.

OpenStudy (imstuck):

You're combining the Law of Sines with Pythagorean's theorem, I think.

OpenStudy (imstuck):



OpenStudy (imstuck):

Can you find the measure of the angle using the Law of Sines?